The measure of the supplement of an angle is 32degrees greater than twice the measure of its complement. Find the measure of the angle.
Answer by jrfrunner(365)

let x=an angle
complement of an angle =90-x, the supplement of an angle =180-x
---
the supplement of an angle is 32degrees greater than twice the measure of its complement
This means supplement=32+2*complement
or 180-x=32+2*(90-x)
180-x=32+180-2x
2x+180-x=32+180 (add 2x to both sides)
2x-x=32 (cancel 180 from both sides)
x=32
-----
check your answer of x=32
go back to the original equation 180-x=32+2*(90-x), subtitute x=32 and see if you get a true statement.
left side : 180-x=180-32=148
right side: 32+2*(90-x)=32+2*(90-32)=32+2*58=32+116=148
so... X=32 makes both sides equal, therefore checked